A plane flies 500 miles with the wind and 350 miles against the wind in the same length of time. If the speed of the wind is 30 mph, what is the speed of the plane in still air?
Answer by HyperBrain(694)

Let x=the speed of the plane in still air
Then x+30=speed of the plane with the wind
x-30=speed of the plane against the wind
---
(1)x+30=500
(2)x-30=350
---
Adding the two equations,
2x=850
x=425 mph